OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring.

OpenStudy (calculusxy):

Are there any quadratic equations that is attached to this question.

OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (calculusxy):

The discriminant, b² - 4ac, is the answer. If b² - 4ac = 0, then the solution to the equation boils down to x = (-b ± 0)/2a = -b/2a, and there is only one solution to the equation, which means the equation cannot be factored by ordinary means. 2x² + 7x + 3 = 0 b² - 4ac = (7)² - 4(2)(3) = 49 - 24 = 25. (yahoo answers)

OpenStudy (anonymous):

they say it has to be a numerical answer

OpenStudy (calculusxy):

Hope that this helps: http://openstudy.com/study#/updates/503ccaebe4b0f52e046a9e51

OpenStudy (goformit100):

First of all @dennisbristol : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct